# Optimal RIS Placement in a Multi-User MISO System with User Randomness

### v1

@@ Abstract
It is well established that the performance of reconfigurable intelligent surface (RIS)-assisted systems critically depends on the optimal placement of the RIS. Previous works consider either simple coverage maximization or simultaneous optimization of the placement of the RIS along with the beamforming and reflection coefficients, most of which assume that the location of the RIS, base station (BS), and users are known. However, in practice, only the spatial variation of user density and obstacle configuration are likely to be known prior to deployment of the system. Thus, we formulate a non-convex problem that optimizes the position of the RIS over the expected minimum signal-to-interference-plus-noise ratio (SINR) of the system with user randomness, assuming that the system employs joint beamforming after deployment. To solve this problem, we propose a recursive coarse-to-fine methodology that constructs a set of candidate locations for RIS placement based on the obstacle configuration and evaluates them over multiple instantiations from the user distribution. The search is recursively refined within the optimal region identified in each stage to determine the final optimal region for RIS deployment. Numerical results are presented to corroborate our findings.
@@ I Introduction
Reconfigurable intelligent surfaces (RIS) have emerged as a promising solution to enhance the performance of a network[1, 2, 3]. An RIS typically operates as a passive reflector that is connected to a smart controller and is capable of adjusting its reflection coefficients based on channel state information (CSI), if available, so that the desired signals are combined constructively and the interfering signals destructively at the receivers[4]. In many cases, its main role in a network is to mitigate blockages caused by buildings and other obstacles[5, 6], a capability that is particularly valuable for higher-carrier-frequency systems, which are known to suffer from severe blockage issues [7, 8]. Given that improving system performance and coverage is the primary utility of an RIS, finding the right place to deploy it in a network is one of the most important and practical problems to address[9, 10].
@@ I-A Related Work
The performance of an RIS-assisted network is enhanced by jointly optimizing the BS beamforming and the reflection coefficients at the RIS (called joint beamforming). However, since a single beamformer and phase configuration cannot effectively serve all possible user combinations from a distribution, this optimization is mostly performed after the system is deployed. [17] presents one of the most cited joint beamforming algorithms that maximizes WSR. Some other strategies are listed in [18, 19, 20]. The machine learning based approaches are not considered because of their dependence on vast training data, which we might not be able to provide in practical scenarios.
@@ II System Model
where $\beta_{br},\beta_{ru_{k}}$ denote the path loss components, $\tilde{\mathbf{H}}_{b{r}}\in\mathbb{C}^{N\times M}$ and $\mathbf{\tilde{h}}_{ru_{k}}\in\mathbb{C}^{N\times 1}$ the unit-variance CSCG components of the links, $\mathbf{\hat{h}_{ru_{k}}}=\mathbf{a}_{M}(\zeta_{k})$ and $\mathbf{\hat{H}_{br}}=\mathbf{a}_{N}(\vartheta)\mathbf{a}_{M}^{H}(\psi)$ the deterministic components of the links. Here, $\mathbf{a}_{i}$ is the steering vector of size $i$ for $i\in\{N,M\}$ and $\vartheta,\psi,\zeta_{k}$ are the angular parameters. $T_{1}$ and $T_{2}$ are the Rician parameters that govern the balance between the steering vector components and the CSCG components. The vectorized reflection matrix of the RIS is given by
$\boldsymbol{\theta}$ $=[\theta_{1},\ldots,\theta_{N}]^{H},$ (4)
where $\theta_{n}=e^{j\varphi_{n}}$ is the phase-shift introduced at the $n^{th}$ element of RIS, for some $1\leq n\leq N$. Let the signal transmitted from the BS be $\mathbf{x}=\sum_{k=1}^{K}\mathbf{w}_{k}s_{k}$, where $\mathbf{w}_{k}\in\mathbb{C}^{M\times 1}$ is the beamforming vector and $s_{k}$ the unit-power transmit information for the $k^{th}$ user. For computational convenience, let us define $\mathbf{H}_{k}=\text{diag}(\mathbf{h}_{{ru}_{k}}^{\mathrm{H}})\mathbf{H}_{br}$. Then, the received signal at user $k$ can be expressed as
$y_{k}=\left(\mathbf{h}_{{bu}_{k}}^{\mathrm{H}}+\boldsymbol{\theta}^{\mathrm{H}}\mathbf{H}_{k}\right)\sum_{k=1}^{K}\mathbf{w}_{k}s_{k}+n_{k},$ (5)
where $n_{k}\sim\mathcal{CN}(0,\sigma^{2})$ is the AWGN at the $k^{th}$ user. The corresponding SINR at user $k$ is
$\gamma_{k}=\frac{\left|\left(\mathbf{h}_{{bu}_{k}}^{\mathrm{H}}+\boldsymbol{\theta}^{\mathrm{H}}\mathbf{H}_{k}\right)\mathbf{w}_{k}\right|^{2}}{\sum_{i\neq k}\left|\left(\mathbf{h}_{{bu}_{k}}^{\mathrm{H}}+\boldsymbol{\theta}^{\mathrm{H}}\mathbf{H}_{k}\right)\mathbf{w}_{i}\right|^{2}+\sigma^{2}}.$ (6)
@@ III Problem Formulation
in which we assume that joint beamforming is optimized using
$\mathbf{P1.2}:\quad(\boldsymbol{\theta},\mathbf{W})=\operatorname*{arg\,max}$ $\sum_{k=1}^{K}\frac{1}{K}\log(1+\gamma_{k})$
@@
where $\mathbf{W}=[\mathbf{w_{1}},\cdots,\mathbf{w_{K}}]\in\mathbb{C}^{M\times K}$ is the accumulated beamformer matrix. The objective in $\mathbf{P1.2}$ is the WSR with equal priority assigned to all users. Eq. (7a) enforces the total transmit power constraint $P_{max}$ for the system. Eq. (7b), on the other hand, defines the feasible set for the RIS reflection coefficients by enforcing unit amplitude at the $N$ elements.
In the proposed problem, we intend to find the optimal location for the RIS that maximizes the expected SINR corresponding to the most affected user (max-min) to enhance the coverage of the system. Let us denote this, $\mathbf{P1.1}$, as the “outer optimization”. It should be noted that the ‘expectation’ is taken over the user distribution. We perform this optimization assuming that joint beamforming is performed after the RIS is deployed at its optimal location. This part of the problem, $\mathbf{P1.2}$, is denoted as the “inner optimization”.
Many researchers consider this multilayered structure of the formulation to be the ‘ideal’ formulation that is most relevant practically. In the next section, we propose an elegant technique to solve this practical problem.
@@ IV Proposed Solution
In this section, we first consider the joint beamforming problem $\mathbf{P1.2}$, then the outer optimization problem $\mathbf{P1.1}$.
@@ IV-A1 Transformed Problem
Using the Lagrangian dual transform technique for closed-form fractional programming [22], problem $\mathbf{P1.2}$ is reduced to a sum-of-ratios problem by introducing a set of auxiliary variables $\boldsymbol{\alpha}=\{\alpha_{k}\}$. Then, the quadratic transform technique [21] is used to decouple the SINR terms (ratios) in the sum-of-ratios problem by introducing another set of auxiliary variables $\boldsymbol{\beta}=\{\beta_{k}\}$, to transform the problem into
$\mathbf{P2.1}:\quad\max_{\boldsymbol{\alpha},\boldsymbol{\beta},\mathbf{W},\boldsymbol{\theta}}$ $f(\boldsymbol{\alpha},\boldsymbol{\beta},\mathbf{W},\boldsymbol{\theta})$
$=$ $\sum_{k=1}^{K}\Big[\frac{1}{K}(\log(1+\alpha_{k})-\alpha_{k})$
$+$ $2\sqrt{\frac{1}{K}(1+\alpha_{k})}\mathcal{R}e\Big\{\beta_{k}^{*}(\mathbf{h}_{{bu}_{k}}^{H}+\boldsymbol{\theta}^{H}\mathbf{H}_{k})\mathbf{w}_{k}\Big\}$
- $|\beta_{k}|^{2}\Big(\sum_{i=1}^{K}|(\mathbf{h}_{{bu}_{k}}^{H}+\boldsymbol{\theta}^{H}\mathbf{H}_{k})\mathbf{w}_{i}|^{2}+\sigma^{2}\Big)\Big]$
@@
$|\theta_{n}|=1,\forall n=1,\cdots,N.$ (8b)
@@ IV-A2 Iterative Optimization
In order to solve this transformed problem, assuming full CSI, we adopt an iterative optimization strategy proposed in [17]. Let $t$ denote the current iteration. The auxiliary variables are updated with their optimal values obtained by setting $\frac{\partial f}{\partial\alpha_{k}}=0$ and $\frac{\partial f}{\partial\beta_{k}}=0$. These optimal values for $\alpha_{k}$ and $\beta_{k}$ are
$\alpha_{k}(t)$ $=\frac{(\zeta_{k}(t-1))^{2}+\zeta_{k}(t-1)\sqrt{(\zeta_{k}(t-1))^{2}+4}}{2},$ (9)
$\beta_{k}(t)$ $=\frac{\mu_{k}(t-1)(\mathbf{h}_{bu_{k}}^{H}+\boldsymbol{\theta}^{H}(t-1)\mathbf{H}_{k})\mathbf{w}_{k}(t-1)}{\sum_{i}|(\mathbf{h}_{bu_{k}}^{H}+\boldsymbol{\theta}^{H}(t-1)\mathbf{H}_{k})\mathbf{w}_{i}(t-1)|^{2}+\sigma^{2}},$ (10)
where $\zeta_{k}(t)=\sqrt{K}\mathcal{R}e\left\{\beta^{*}_{k}(t)(\mathbf{h}_{bu_{k}}^{H}+\boldsymbol{\theta}^{H}(t)\mathbf{H}_{k})\mathbf{w}_{k}(t)\right\}$ and $\mu_{k}(t)=\sqrt{\frac{1}{K}(1+\alpha_{k}(t))}$. Here, $\mathcal{R}e\{.\}$ represents the real part. The update rule for $\mathbf{W}$ is given as
$\mathbf{w}_{k}(t)$ $=\frac{1}{L(t)-2\lambda(t)}(L(t)\hat{\mathbf{w}}_{k}(t-1)-\mathbf{g}_{k}(t)),$ (11)
[…]
where $\hat{\mathbf{w}}_{k}(t)=\mathbf{w}_{k}(t)+\epsilon(\mathbf{w}_{k}(t)-\mathbf{w}_{k}(t-1))$ is the extrapolated beamformer for some $\epsilon\geq 0$ and $\mathbf{g}_{k}(t)=-\frac{\partial f_{1}}{\partial\mathbf{w_{k}}}$ in which $f_{1}(\mathbf{W})=f(\boldsymbol{\alpha}(t-1),\boldsymbol{\beta}(t-1),\mathbf{W},\boldsymbol{\theta}(t-1))$. Also, $L(t)$ is set to the Lipschitz constant of $\mathbf{g}_{k}(t)$.
Finally, the update rule for phase coefficients $\boldsymbol{\varphi}=[\varphi_{1},\cdots,\varphi_{N}]^{T}$ (recall $\theta_{n}=e^{j\varphi_{n}}$) is derived based on the second-order Taylor series expansion of $f_{2}(\boldsymbol{\varphi})=f(\boldsymbol{\alpha}(t-1),\boldsymbol{\beta}(t-1),\mathbf{W}(t-1),e^{j\boldsymbol{\varphi}})$ to be of the form
$\boldsymbol{\varphi}(t)\leftarrow\boldsymbol{\varphi}(t-1)-\frac{\nabla f_{2}(\boldsymbol{\varphi}(t-1))}{\kappa},$ (13)
where optimal $\kappa$ is obtained using the strategy listed in [17].

### v2

# Optimal RIS Placement in Multi-User MISO Systems with User Randomness

@@ Abstract
It is well established that the performance of reconfigurable intelligent surface (RIS)-assisted systems critically depends on the optimal placement of the RIS. Previous works consider either simple coverage maximization or simultaneous optimization of the placement of the RIS along with the beamforming and reflection coefficients, most of which assume that the location of the RIS, base station (BS), and users are known. However, in practice, only the spatial variation of user density and obstacle configuration are likely to be known prior to deployment of the system. Thus, we formulate a non-convex problem that optimizes the position of the RIS over the expected minimum signal-to-interference-plus-noise ratio (SINR) of the system with user randomness, assuming that the system employs joint beamforming after deployment. To solve this problem, we propose a recursive coarse-to-fine methodology that constructs a set of candidate locations for RIS placement based on the obstacle configuration and evaluates them over multiple instantiations from the user distribution. The search is recursively refined within the optimal region identified in each stage to determine the final optimal region for RIS deployment. Detailed numerical results are presented to corroborate our findings.
@@ I Introduction
Reconfigurable intelligent surfaces (RIS) have emerged as a promising solution to enhance the performance of a network[1, 2, 3]. An RIS embedded with a large number of small reconfigurable elements typically operates as a passive reflector that is connected to a smart controller and is capable of adjusting its reflection coefficients based on channel state information (CSI), if available, so that the desired signals are combined constructively and the interfering signals destructively at the receivers[4, 5]. In many cases, its main role in a network is to mitigate blockages caused by buildings and other obstacles[6, 7], a capability that is particularly valuable for higher-carrier-frequency systems, which are known to suffer from severe blockage issues [8, 9]. Given that improving system performance and coverage is the primary utility of an RIS, finding the right place to deploy it in a network is one of the most important and practical problems to address[10, 11].
@@ I-A Related Work
The performance of an RIS-assisted network is enhanced by jointly optimizing the BS beamforming and the reflection coefficients at the RIS (called joint beamforming). However, since a single beamformer and phase configuration cannot effectively serve all possible user combinations from a distribution, this optimization is mostly performed after the system is deployed. [18] presents one of the most accepted joint beamforming algorithms that maximizes WSR. Some other strategies are listed in [19, 20, 22]. The machine learning based approaches are not considered because of their dependence on vast training data, which we might not be able to provide in practical scenarios.
@@ II System Model
where $\beta_{br}$ and $\beta_{ru_{k}}$ denote the path loss components of the BS-RIS and RIS-User $k$ links, respectively. $\tilde{\mathbf{H}}_{b{r}}\in\mathbb{C}^{N\times M}$ and $\mathbf{\tilde{h}}_{ru_{k}}\in\mathbb{C}^{N\times 1}$ denote the unit-variance CSCG components of the corresponding links, and $\mathbf{\hat{H}}_{br}=\mathbf{a}_{N}(\vartheta)\mathbf{a}_{M}^{H}(\psi)$ and $\mathbf{\hat{h}}_{ru_{k}}=\mathbf{a}_{N}(\zeta_{k})$ denote the deterministic components of the links. Here, $\mathbf{a}_{i}$ is the steering vector of size $i$ for $i\in\{N,M\}$ and $\vartheta,\psi,\zeta_{k}$ are the angular parameters. $T_{1}$ and $T_{2}$ are the Rician parameters that govern the balance between the steering vector components and the CSCG components. The reflection matrix of the RIS is given by
$\boldsymbol{\Theta}$ $=\text{diag}(\boldsymbol{\theta}),$ (4)
$\text{where }\boldsymbol{\theta}$ $=[\theta_{1},\ldots,\theta_{N}]^{T},$ (5)
in which, $\theta_{n}=e^{j\varphi_{n}}$ is the phase-shift introduced at the $n^{th}$ element of RIS, for all $1\leq n\leq N$. Let the combined signal transmitted from the BS be $\mathbf{x}=\sum_{k=1}^{K}\mathbf{w}_{k}s_{k}$, where $\mathbf{w}_{k}\in\mathbb{C}^{M\times 1}$ is the transmit beamformer and $s_{k}$ the unit-power transmit symbol corresponding to the $k^{th}$ user. The signal received at user $k$ can be expressed as
$y_{k}$ $=\mathbf{h}_{{bu}_{k}}^{\mathrm{H}}\mathbf{x}+\mathbf{h}_{ru_{k}}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{H}_{br}\mathbf{x}+n_{k}$
$=\left(\mathbf{h}_{{bu}_{k}}^{\mathrm{H}}+\mathbf{h}_{ru_{k}}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{H}_{br}\right)\sum_{k=1}^{K}\mathbf{w}_{k}s_{k}+n_{k},$ (6)
where $n_{k}\sim\mathcal{CN}(0,\sigma^{2})$ is the AWGN at the $k^{th}$ user. For computational convenience, the indirect links are redefined using $\mathbf{H}_{r,k}=\text{diag}(\mathbf{h}_{{ru}_{k}}^{\mathrm{H}})\mathbf{H}_{br}$. Thus, the revised expression for the received signal at user $k$ is given by
$y_{k}=\left(\mathbf{h}_{{bu}_{k}}^{\mathrm{H}}+\boldsymbol{\theta}^{\mathrm{H}}\mathbf{H}_{r,k}\right)\sum_{k=1}^{K}\mathbf{w}_{k}s_{k}+n_{k},$ (7)
Let $\mathbf{h}_{k}\triangleq(\mathbf{h}_{{bu}_{k}}^{H}+\boldsymbol{\theta}^{H}\mathbf{H}_{r,k})$ denote the end-to-end channel for user $k$. Thus, the SINR at user $k$ is
$\gamma_{k}=\frac{\left|\mathbf{h}_{k}\mathbf{w}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}\mathbf{w}_{i}\right|^{2}+\sigma^{2}}.$ (8)
@@ III Problem Formulation
$=\max_{\mathbf{o}_{r}}$ $\quad\mathbb{E}_{\mathbf{o}_{u}}\left[\min\frac{\left|\mathbf{h}_{k}\mathbf{w}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}\mathbf{w}_{i}\right|^{2}+\sigma^{2}}\right]$
in which we assume that $\boldsymbol{\theta}$ in $\mathbf{h}_{k}$ and the accumulated beamformer matrix $\mathbf{W}=[\mathbf{w_{1}},\cdots,\mathbf{w_{K}}]\in\mathbb{C}^{M\times K}$ are optimized using
$\mathbf{P1.2}:\quad(\boldsymbol{\theta},\mathbf{W})=\arg\max_{\boldsymbol{\theta,\mathbf{W}}}$ $\sum_{k=1}^{K}\frac{1}{K}\log(1+\gamma_{k})$
@@
The objective function in $\mathbf{P1.2}$ is the average rate or WSR with equal priority assigned to all users. Eq. (9a) enforces the total transmit power constraint $P_{max}$ for the system. Eq. (9b), on the other hand, defines the feasible set for the RIS reflection coefficients by enforcing unit amplitude at all $N$ elements.
In problem $\mathbf{P1.1}$, we intend to find that RIS placement that maximizes the expected SINR of the most-affected user among the $K$ users (that has the least SINR) to enhance the coverage of the system. This is called ‘max-min’ solution which is the most “fair” one. This problem of finding the most fair solution for the placement of RIS is denoted as the “outer optimization”. It should be noted that this ‘expectation’ of SINR is taken over the user distribution. We perform this outer optimization assuming that joint beamforming is performed after the RIS is deployed at its optimal location. This part of the problem, $\mathbf{P1.2}$, is denoted as the “inner optimization”.
Many researchers consider this multilayered structure of the formulation to be the ‘ideal’ one that is closest to the practical scenario. However, because of the difficulty in directly solving it, existing works like [13, 14, 12] simplify the formulation in various ways. In the next section, we propose an elegant technique to directly tackle the practical hierarchical formulation.
@@ IV Proposed Solution
In order to solve $\mathbf{P1.1}$, we need the optimal beamformer and phases from $\mathbf{P1.2}$. Therefore, in this section, we first consider solving the joint beamforming problem $\mathbf{P1.2}$ and then the outer optimization problem $\mathbf{P1.1}$.
@@ IV-A1 Transformed Problem
Using the Lagrangian dual transform [24], $\mathbf{P1.2}$ is reduced to a sum-of-ratios problem by introducing a set of auxiliary variables $\boldsymbol{\alpha}=\{\alpha_{1},\dots,\alpha_{K}\}$
$\mathbf{P2.1}:\quad\max_{\boldsymbol{\alpha},\mathbf{W},\boldsymbol{\theta}}$ $\sum_{k=1}^{K}\Big[\frac{1}{K}(\log(1+\alpha_{k})-\alpha_{k}+\frac{(1+\alpha_{k})\gamma_{k}}{1+\gamma_{k}}\Big]$
@@
$|\theta_{n}|=1,\quad\forall n=1,\cdots,N,$ (10b)
$\alpha_{k}\geq 0,\quad\forall k=1,\cdots,K.$ (10c)
Clearly, the problem is equivalent to $\mathbf{P1.2}$ as long as $\alpha_{k}=\gamma_{k}$. Then, the quadratic transform [23] is applied to decouple the ratios (SINR) in $\mathbf{P2.1}$ by introducing another set of auxiliary variables $\boldsymbol{\beta}=\{\beta_{1},\dots,\beta_{K}\}$
$\mathbf{P2.2}:\quad\max_{\boldsymbol{\alpha},\boldsymbol{\beta},\mathbf{W},\boldsymbol{\theta}}$ $\sum_{k=1}^{K}\Big[\frac{1}{K}(\log(1+\alpha_{k})-\alpha_{k})$
$+$ $2\sqrt{\frac{1}{K}(1+\alpha_{k})}\mathcal{R}e\Big\{\beta_{k}^{*}\mathbf{h}_{k}\mathbf{w}_{k}\Big\}$
- $|\beta_{k}|^{2}\Big(\sum_{i=1}^{K}|\mathbf{h}_{k}\mathbf{w}_{i}|^{2}+\sigma^{2}\Big)\Big]$
$s.t.\quad$ $\sum_{k=1}^{K}\|\mathbf{w}_{k}\|_{2}^{2}\leq P_{max},$ (11a)
$|\theta_{n}|=1,\forall n=1,\cdots,N,$ (11b)
$\alpha_{k}\geq 0,\quad\forall k=1,\cdots,K.$ (11c)
@@ IV-A2 Iterative Optimization
In order to solve $\mathbf{P2.2}$, we adopt an alternative optimization (AO) strategy. The auxiliary variables are updated with their optimal values [21]
$\alpha_{k}$ $=\gamma_{k},$ (12)
$\beta_{k}$ $=\frac{\sqrt{\frac{1}{K}(1+\alpha_{k})}\,\mathbf{h}_{k}\mathbf{w}_{k}}{\sum_{i}|\mathbf{h}_{k}\mathbf{w}_{i}|^{2}+\sigma^{2}}.$ (13)
Looking at the well-studied optimal transmit beamforming problem with constant phases, one can derive the update rule for $\mathbf{W}$ to be the WMMSE solution which is given by
$\mathbf{w}_{k}$ $=\sqrt{(1+\alpha_{k})}\beta_{k}\Big(\lambda^{\text{opt}}\mathbf{I}_{M}+\sum_{i=1}^{K}|\beta_{i}|^{2}\mathbf{h}_{i}\mathbf{h}_{i}^{H}\Big)^{-1}\mathbf{h}_{k},$ (14)
$\lambda^{\text{opt}}$ $=\min\left\{\lambda^{\text{opt}}\geq 0\mid\sum_{i=1}^{K}|\mathbf{w}_{k}|^{2}\leq P_{max}\right\},$ (15)
where $\lambda^{\text{opt}}$ is obtained by solving Eq. (15) using a bisection search over a finite interval that starts at 0 (upto $10^{9}$, let’s say). Finally, one can see that the phase dependent terms in the objective from $\mathbf{P2.2}$ can be shown [18] to be of the form
$\mathbf{P2.3}\quad$ $\min_{\boldsymbol{\varphi}}f_{3}(\boldsymbol{\varphi})=(e^{j\boldsymbol{\varphi}})^{H}\boldsymbol{U}e^{j\boldsymbol{\varphi}}-2\mathcal{R}e\{\boldsymbol{v}^{H}e^{j\boldsymbol{\varphi}}\}$ (16)
$\text{where }\quad\boldsymbol{U}$ $=\sum_{k=1}^{K}|\beta_{k}|^{2}\sum_{i=1}^{K}\mathbf{H}_{r,k}\mathbf{w}_{i}(\mathbf{H}_{r,k}\mathbf{w}_{i})^{H},$ (17)
$\boldsymbol{v}$ $=\sum_{k=1}^{K}\Big(\sqrt{\frac{1}{K}(1+\alpha_{k})}\beta^{*}_{k}(\mathbf{H}_{r,k}\mathbf{w}_{i})$
$-|\beta_{k}|^{2}\sum_{i=1}^{K}(\mathbf{h}^{H}_{bu,k}\mathbf{w}_{i})^{*}\mathbf{H}_{r,k}\mathbf{w}_{i}\Big),$ (18)
Hence, to optimize phases with constant transmit beamforming, we adopt the gradient projection algorithm
$\boldsymbol{\varphi}\leftarrow\left(\boldsymbol{\varphi}-\rho\nabla f_{3}(\boldsymbol{\varphi})\right)^{+},$ (19)
where $(.)^{+}$ denotes the projection of the enclosed set of phases into the region $|\theta_{n}|=1,\forall n=1,\cdots,N$. Also, $\rho$ is dynamically updated in each iteration for faster convergence.
We alternatively optimize $(\boldsymbol{\alpha},\boldsymbol{\beta},\mathbf{W},\boldsymbol{\theta})$ using Eq.  (12), (13), (14) and (19) iteratively to solve the inner optimization problem.